OpenStudy (anonymous):

A 8.5 L sample is collected at 45°C. What will the volume be if the temperature is raised to 70°C? A. 9.2 L B. 13.2 L C. 5.5 L D. 7.9 L

OpenStudy (zpupster):

V1/T1=V2/T2 V1=8.5 T1=45 V2=x T2=70 8.5(70)/45=x

OpenStudy (anonymous):

I got D as my answer.

OpenStudy (zpupster):

do what i showed in calculator

OpenStudy (anonymous):

Oh! Okay, my bad, its B.